# Cross-D Conv: Cross-Dimensional Transferable Knowledge Base via Fourier Shifting Operation

### v3

@@ II Methodology
Rotation Parameter Representation. Let $x\in\mathbb{R}^{4\times H\times W}$ be the input to the secondary network, and let $f(x)\in\mathbb{R}^{4\times(HW)}$ be the corresponding feature map it produces. We aggregate these features into a single 4D rotation parameter vector $r$ using a softmax-weighted sum:
@@ III Experiments
Datasets For pretraining, we utilized the ImageNet dataset, which comprises 1,000 classes of natural images, alongside its biomedical counterpart, RadImageNet. RadImageNet encompasses approximately 1.35 million annotated medical images across CT, MRI, and ultrasound modalities, categorized into 165 classes spanning 11 anatomical regions.
@@
Baseline Model We selected the ResNet-18 architecture due to its modern features, such as skip connections, varying kernel sizes, and batch normalization, it can replace any 2D convolution operation. ResNet-18 + 3D: This is a 3D extension of the ResNet-18 architecture that can process volumetric data directly. ResNet-18 + ACS Conv [2]: Axial-Coronal-Sagittal Convolutions split 2D kernels into three parts along the channel dimension and apply them separately to the three orthogonal views of a 3D volume. This approach allows the use of pre-trained 2D weights while enabling 3D representation learning. ACS-Conv networks can be initialized with both ImageNet and RadImageNet [3] priors.
@@ V Conclusion
Blood [16] (Microscope) 17,092 (224, 224, 3) 0 – 255 8 Multi-class
Brain [7] (MRI) 1,600 (224, 224, 3) 0 – 255 23 Multi-class
Brain Tumor [6] (MRI) 3,064 (224, 224, 3) 0 – 255 3 Multi-class
Breast Cancer [9] (US) 1,875 (224, 224, 3) 0 – 255 2 Binary
Breast [8] (US) 780 (224, 224, 1) 0 – 255 2 Binary
Derma [18] (Dermatology) 10,015 (224, 224, 3) 0 – 255 7 Multi-class
OrganC [4] (CT) 23,582 (224, 224, 1) 0 – 255 11 Multi-class
OrganS [5] (CT) 25,211 (224, 224, 1) 0 – 255 11 Multi-class
Pneumonia [17] (XR) 5,856 (224, 224, 1) 0 – 255 2 Binary
@@
Brain Tumor: With 3,064 MRI images divided into 3 classes, this dataset focuses on the detection and classification of brain tumors. It is instrumental in assessing the model’s sensitivity and specificity in medical diagnosis applications.
@@
OrganS: With 25,211 CT images labeled into 11 classes, this dataset is similar to OrganC but focuses on segmentation tasks, providing a basis for assessing models’ abilities in delineating organ boundaries.
@@
BraTS21: This dataset includes 585 MRI scans with binary labels indicating the presence or absence of brain tumors. The volumetric nature of the data (96x96x96 voxels) challenges models to accurately segment and classify tumor regions.
@@
Synapse: Comprising 1,759 microscope volumes with binary labels, this dataset is used for evaluating models in segmenting and classifying synapses in neural imaging.
Vessel: This dataset contains 1,908 MRA volumes with binary labels, focusing on vessel segmentation tasks, essential for evaluating models in vascular imaging applications.

### v4

@@ II Methodology
Rotation Parameter Representation. Let $x\in\mathbb{R}^{C\times H\times W}$ be the input to the secondary network, and let $f(x)\in\mathbb{R}^{4\times(HW)}$ be the corresponding feature map it produces. We aggregate these features into a single 4D rotation parameter vector $r$ using a softmax-weighted sum:
@@ III Experiments
Datasets For pretraining, we utilized the ImageNet dataset, which comprises 1,000 classes of natural images, alongside its biomedical counterpart, RadImageNet which encompasses approximately 1.35 million annotated medical images across CT, MRI, and ultrasound modalities, categorized into 165 classes spanning 11 anatomical regions.
@@
Baseline Model We selected the ResNet-18 architecture due to its modern features, such as skip connections, varying kernel sizes, and batch normalization. But typically, Cross-D Conv can replace any 2D convolution operation. ResNet-18 + 3D: This is a 3D extension of the ResNet-18 architecture that can process volumetric data directly. ResNet-18 + ACS Conv [2]: Axial-Coronal-Sagittal Convolutions split 2D kernels into three parts along the channel dimension and apply them separately to the three orthogonal views of a 3D volume. This approach allows the use of pre-trained 2D weights while enabling 3D representation learning. ACS-Conv networks can be initialized with both ImageNet and RadImageNet [3] priors.
@@ V Conclusion
Blood [16] Microscope 17,092 (224, 224, 3) 0 – 255 8 Multi-class
Brain [7] MRI 1,600 (224, 224, 3) 0 – 255 23 Multi-class
Brain Tumor [6] MRI 3,064 (224, 224, 3) 0 – 255 3 Multi-class
Breast Cancer [9] US 1,875 (224, 224, 3) 0 – 255 2 Binary
Breast [8] US 780 (224, 224, 1) 0 – 255 2 Binary
Derma [18] Dermatology 10,015 (224, 224, 3) 0 – 255 7 Multi-class
OrganC [4] CT 23,582 (224, 224, 1) 0 – 255 11 Multi-class
OrganS [5] CT 25,211 (224, 224, 1) 0 – 255 11 Multi-class
Pneumonia [17] XR 5,856 (224, 224, 1) 0 – 255 2 Binary
@@
Brain Tumor: With 3,064 MRI images divided into 3 classes, this dataset focuses on classification of brain tumors. It is instrumental in assessing the model’s sensitivity and specificity in medical diagnosis applications.
@@
OrganS: With 25,211 CT images labeled into 11 classes, this dataset is similar to OrganC.
@@
BraTS21: This dataset includes 585 MRI scans with binary labels indicating the presence or absence of brain tumors. The volumetric nature of the data (96x96x96 voxels) challenges models to accurately classify tumor regions.
@@
Synapse: Comprising 1,759 microscope volumes with binary labels, this dataset is used for evaluating models in classifying synapses in neural imaging.
Vessel: This dataset contains 1,908 MRA volumes with binary labels, focusing on vessel classification tasks, essential for evaluating models in vascular imaging applications.